OpenStudy (anonymous):

Simplify. 5 • 2 • (–4) + 3 • 6

OpenStudy (anonymous):

U would do all the multiplication first then add the 3.

OpenStudy (anonymous):

\(\Large 5\times2\times(-4)+3\times6=-40+18=-22\)